3y^3-2y[4y-y(y-3)]-[2y(y+1)-3y(y^2-1)]

Answer by Boreal(15235)

3y^3-2y[4y-y(y-3)]-[2y(y+1)-3y(y^2-1)]=
3y^3-2y(4y-y^2+3y)-(2y^2+2y-3y^3+3y)=
3y^3-8y^2+2y^3-6y^2-2y^2-2y+3y^3-3y=
8y^3-16y^2-5y
=y(8y^2-16y-5)